# Free Majorana Fermion Meets Gauged Ising Conformal Field Theory on the Fuzzy Sphere

### v2

@@ 4 Critical Theories
In this section, we discuss the critical theories that describe the transition between the TSC and fIQH and bPf phases.
The transition from the fIQH to the TSC phase is described by a single free Majorana fermion. One way to see this is that the chiral central charge $c_{-}$ increases by $1/2$ across the transition. This transition is captured by the Lagrangian
@@
where $\chi$ is a (two-component) Majorana fermion, $\bar{\chi}=\chi^{\mathrm{T}}\epsilon$, $\epsilon=i\sigma^{2}$, and $\not{\partial}=\sigma^{\mu}\partial_{\mu}$. The gravitational Chern-Simons term is defined through extending into a 4D manifold $X$ with the original 3D manifold its boundary $M=\partial X$ [77]
@@
where $R$ is the curvature 2-form. A background $\text{CS}_{g}$ level-1 comes from the background integer quantum Hall state has been included. Adding a positive/negative mass gaps out the Majorana fermion and adds an extra gravitational Chern-Simons response of level $\pm 1/4$.
@@
They recover the chiral central charge of the fIQH and TSC phases. This transition has another conjectural dual description of $\mathrm{SO}(3)_{1}$ coupled to a single free scalar [78, 79]. The condensation of the scalar Higgses the gauge theory to $\mathrm{SO}(2)_{1}$ describing the fIQH phase, and gapping the scalar results in $\mathrm{SO}(3)_{1}$ corresponding to the TSC phase.
The transition from the TSC to the bPf phase falls into a gauged Ising (Ising) universality, i. e. Ising type transition coupled to a $\mathbb{Z}_{2}$ gauge field, or the confinement transition of a $\mathbb{Z}_{2}$ gauge field. One way to understand it is to recognise that bPf can be described by TSC coupled to a $\mathbb{Z}_{2}$ gauge field, so its transition to TSC can be realised by Higgsing the $\mathbb{Z}_{2}$ gauge field. We can also use the language of CS theory to understand the emergence of gauged Ising transition. The TSC and the bPf phase can be described by $\mathrm{SO}(3)_{1}\cong\mathrm{SU}(2)_{2}/\mathbb{Z}_{2}$ and $\mathrm{SU}(2)_{2}$ respectively, and the difference is only a $\mathbb{Z}_{2}$ gauge field. So Higgsing the $\mathbb{Z}_{2}$ centre of the $\mathrm{SU}(2)$ gauge group drives a transition from the bPf phase to the TSC phase. In Appendix B, we discuss the Lagrangian of this confinement transition. This transition can be equivalently considered in terms of anyon condensation. In the bPf phase, although the $\psi$ anyon is fermionic, combined with the local fermion $f$, $f^{\dagger}\psi$ is bosonic with statistical angle $\theta_{f^{\dagger}\psi}=0$. Across the transition, this composite particle $\langle f^{\dagger}\psi\rangle\neq 0$ is condensed. Due to the braiding of the anyons $\psi$ and $\sigma$, $\sigma$ will also be confined, resulting in the $f$-wave TSC with trivial topological order. Since this transition only changes the neutral sector, the chiral central charge $c_{-}=3/2$ remains the same across the transition.
It is also worth noting that the whole phase diagram, including the three gapped phases and two critical points, can be captured by a single Lagrangian, viz. the adjoint QCD of $\mathrm{SU}(2)_{1}$ Chern-Simons theory coupled to a fermion in the adjoint representation of the gauge group [65]. This theory has no global symmetry. When the fermion is gapped, it contributes a Chern-Simons level $\pm 1$ depending on the sign of the mass. This drives a transition between $\mathrm{SU}(2)_{2}$ topological order and a trivial topological order of $\mathrm{SU}(2)_{0}$. This transition is dual to $\mathrm{O}(1)_{-3}$ coupled to a real scalar through the level-rank duality, which is equivalent to the Ising universality. On the other hand, the adjoint QCD has a manifest $\mathscr{N}=1$ supersymmetry for a specific value of mass. This point falls in the phase with trivial topological order. The supersymmetry is spontaneously broken, resulting in a gapless Goldstino mode: a massless Majorana fermion. The supersymmetry is not manifest in our setup, but the Goldstino mode corresponds to the Majorana-type transition in our phase diagram.
@@
Figure 4: The rescaled energy spectrum at the conformal points (a) $t=0.3,\mu=0.0$ described by the free Majorana fermion CFT and (b) $t=0.1,\mu=2.25$ described by the Ising CFT. We identify some multiplets of the lowest primaries in the spectra and the bars denote their expected value in the CFT. The spectra are calibrated by $\Delta_{T}=3$. In the calculation, we fix $N_{mf}=12$.
At the fIQH-TSC transition, we observe that most of the operators are close to integer level and they can be classified into multiplets of the lowest primary operators in the free Majorana fermion CFT (Figure 4a), viz. the fermion $\chi$ at $\Delta=1,l=1/2$, the relevant singlet $\bar{\chi}\chi=\epsilon_{ab}\chi^{a}\chi^{\beta}$ at $\Delta=2,l=0$, the conserved stress tensor $T^{\mu\nu}=\tfrac{i}{2}\bar{\chi}_{a}(\sigma^{(\mu})^{a}{}_{b}\partial^{\nu)}\chi^{b}$ at $\Delta=3,l=2$, $\bar{\chi}\chi\partial\chi$ at $\Delta=4,l=3/2$, etc. We highlight that spectrum contains $l\in\mathbb{Z}+\tfrac{1}{2}$ states corresponding to fermionic operators. The free Majorana fermion CFT will be analysed in more detail in Section 5.
At the TSC-bPf transition, although there are many non-conformal gapped excitations at relatively low energy at small system size, the low lying states for each integer $l\leq 2$ can be identified as operators in the Ising CFT (Figure 4b). The Ising CFT consists all the $\mathbb{Z}_{2}$-even operators in the Ising CFT. We can identify $\epsilon$ at $\Delta=1.412$ and its several descendants, the stress tensor $T^{\mu\nu}$ and $\epsilon^{\prime}$ at $\Delta=3.830$ in the spectrum. The Ising CFT will be discussed in more detail in Section 6.
5 Free Majorana Fermion CFT
In this section, we focus on the free Majorana fermion CFT on the fuzzy sphere. In this CFT, the Majorana fermion $\chi$ has scaling dimension $\Delta=1$ and $\mathrm{SU}(2)$ spin $l=1/2$. The spectrum of scaling operator, including the primaries and descendants, can be determined with the help of harmonic oscillators [46]. Here we list the scaling dimension $\Delta$ and spin $l$ of the low lying operators:
@@ 5 Free Majorana Fermion CFT
Figure 5: The spectrum with $\Delta\lesssim 4$ for various systems sizes. The scaling dimensions are calibrated by the stress tensor with $\Delta_{T}=3$. The grey bars denote the theoretical value for comparison. In the calculation, we fix $t=0.3$ and $\mu=0$. The verticle gridlines with interval $0.5$ separates differnt spins, which are either integers or half-integers.
We find that the low lying spectrum agrees well with the theory (Figure 5). All the states up to $\Delta=4$ are CFT states, and all the theoretically expected operators appear in the spectrum close to integer scaling dimension. The discrepancy is within $2.5\%$ for all the $\Delta\leq 3$ operators and most $\Delta=4$ operators for our maximal accessible size $N_{mf}=13$. This precision is remarkable with almost no fine-tuning at such small size even compared with similar fuzzy sphere model (e. g. the free real scalar [46]). With increasing system size, the scaling dimension gradually approaches to expected integer value (Figure 6a–d). We qualitatively evaluate this through the cost function consisting of all the operators with $\Delta\leq\Delta_{c}=2,3,4$. We find these cost functions all scales to zero as $Q\sim N_{mf}^{-\alpha/2}$ where $\alpha=3.4,3.6,3.2$ respectively for $\Delta_{c}=2,3,4$ (Figure 6e).
Figure 6: (a–d) The finite size scaling of scaling dimensions of various operators. The grey lines denote the theoretical value for comparison. (e) The finite size scaling of cost function consisting of operators with $\Delta\leq\Delta_{c}=2,3,4$. In the calculation, we fix $t=0.3$ and $\mu=0$.
@@
$f^{\dagger}(\mathbf{r})b(\mathbf{r})$ $=\alpha\chi^{(1/2)}(\mathbf{r})+\dots$
$b^{\dagger}(\mathbf{r})f(\mathbf{r})$ $=\bar{\alpha}\chi^{(-1/2)}(\mathbf{r})+\dots,$ (8)
@@
$\langle\chi^{a}(x_{1})\bar{\chi}_{b}(x_{2})\rangle=\frac{(\sigma_{\mu})^{a}{}_{b}x_{12}^{\mu}}{(x_{12}^{2})^{\Delta_{\chi}+1/2}},$ (9)
where $\Delta_{\chi}=1$, $x_{1,2}\in\mathbb{R}^{3}$, $x_{12}=x_{1}-x_{2}$, $a,b=\pm$ are spinor indices, and $\sigma$ are the Pauli matrices. Taking the polarisations $\chi^{(s)}(\mathbf{r})$ on the sphere, their two-point function
$\langle 0|\chi^{\dagger\,(s_{1})}(\mathbf{r}_{1})\chi^{(s_{2})}(\mathbf{r}_{2})|0\rangle=\frac{2(-1)^{s_{1}}\delta_{s_{1}s_{2}}}{R^{2\Delta_{\chi}}(2\sin\frac{1}{2}\gamma_{12})^{2\Delta_{\chi}}}$ (10)
has the same form as scalars on the sphere, where $\gamma_{12}$ is the angular distance between the two points.
Figure 7: (a) The dimensionless correlation function Eq. (11) at diffrent system sizes. The grey dashed line denotes $(2\sin\gamma_{12}/2)^{-2}$ required by conformal symmetry. (b) The scaling dimension $\Delta_{\chi}$ extractd from the corrlation function through different ways in Eqs. (12) as a function of system size. In the calculation, we fix $t=0.3$ and $\mu=0$.
@@
$C_{\chi}^{\ast}(\gamma_{12})=\tfrac{1}{2i}R^{2\Delta_{\chi}}\langle 0|\chi^{\dagger\,(1/2)}(\mathbf{r}_{1})\chi^{(1/2)}(\mathbf{r}_{2})|0\rangle.$ (11)
whose exact expression is given in Eq. (42), and plot the results in Figure 7a. As the system size increases, the numerical result gradually approaches its theoretical value $(2\sin\gamma_{12}/2)^{-2\Delta_{\chi}}$. From this correlation function, the scaling dimension $\Delta_{\chi}$ can be extracted in several ways
$\left.C_{\chi}^{\ast}(\gamma_{12})\right|_{\gamma_{12}=\pi}$ $=2^{-2\Delta_{\chi}}$ (12a)
$\left.\frac{(C^{\ast}_{\chi})^{\prime\prime}(\gamma_{12})}{C^{\ast}_{\chi}(\gamma_{12})}\right|_{\gamma_{12}=\pi}$ $=\frac{\Delta_{\chi}}{2}$ (12b)
$\int_{-1}^{1}\mathrm{d}\cos\gamma_{12}\,\sin\frac{\gamma_{12}}{2}C^{\ast}_{\chi}(\gamma_{12})$ $=\frac{2^{2(1-\Delta_{\chi})}}{3-2\Delta_{\chi}}.$ (12c)
@@
We also note that under this expression of $\chi$, the pair conversion term $H_{t}$ is proportional to
$H_{t}\propto\int\mathrm{d}^{2}\mathbf{r}\,i\bar{\chi}_{a}(\mathbf{r})(\sigma_{\mu})^{a}{}_{b}\partial^{\mu}\chi^{b}(\mathbf{r})$ (13)
which coincides with the kinetic term in the Hamiltonian of free Majorana fermion.
@@ 6 Gauged Ising CFT
Numerically, we can indeed identify $\epsilon,\epsilon^{\prime},T^{\mu\nu}$ and their descendants up to $\Delta<5$ and $l\leq 2$ in the spectrum (Figure 8a). In the calculation, we fix $t=0.1$ with a relatively optimal conformal symmetry, and choose $\mu$ for each system size by optimising a cost function consisting of $\epsilon,\partial^{\mu}\epsilon,\partial^{\mu}\partial^{\nu}\epsilon,\Box\epsilon$ and $T^{\mu\nu}$. The optimal $\mu$ for each size is listed in Table 1 (the meaning of the right panel will be explained shortly). We note that although there are non-conformal states due to the lowest gapped excitations in the proximate bosonic Pfaffian phase, the low energy spectrum with $l\leq 2$ show relatively clear conformal signature. At large enough system size, the low energy specctrum should be dominated by CFT states and these gapped excitations will move higher, but at the sizes available, these excitations appear at relatively low energy at large $l$. As the bosonic Pfaffian state is only compatible with even $N_{mf}$, and the optimal conformal point drifts with system size, we leave the finite size scaling analysis to future work.
@@
Figure 8: The spectrum of the identified CFT states with $\Delta<5$ for various systems sizes. The grey bars denote the theoretical value for comparison. (a) The bulk spectrum at even $N_{mf}$ corresponding to the $\mathbb{Z}_{2}$-even part of the Ising CFT spectrum. The scaling dimensions are calibrated by the stress tensor with $\Delta_{T}=3$. (b) The Wilson line endpoint spectrum at odd $N_{mf}$ corresponding to the $\mathbb{Z}_{2}$-odd part of the Ising CFT spectrum. The scaling dimensions are calibrated by setting $\Delta_{\sigma}=0.518$ and $\Delta_{\partial^{\mu}\sigma}=1.518$. In the calculation, we fix $t=0.1$ and choose $\mu$ as in Table 1.
@@
On the other hand, the $\mathbb{Z}_{2}$-odd operators can be made gauge invariant by attaching a Wilson line (of $\mathbb{Z}_{2}$ gauge field) to it. In other words, the $\mathbb{Z}_{2}$-odd operators become endpoint operators of a line defect in the gauged Ising CFT. Although the defect endpoint operators arise generically in conformal line defects and have been discussed in various contexts [80, 81, 30, 82], in the setting of $\mathbb{Z}_{2}$ (or any discrete) gauge theory, the Wilson line defect is not only conformal but also topological. In the bPf-TSC transition, we can straightforwardly realise the topological Wilson line on fuzzy sphere by leveraging the even-odd effect of $N_{mf}$.
@@
We first note that the bosonic Pfaffian state is sensitive to the parity of particle number, and is only compatible with even $N_{mf}$. The choice of the odd $N_{mf}$, on the other hand, is equivalent to the insertion of a gauge charge at the centre of the sphere as the endpoint of a Wilson line defect. The reason is that although the $\mathrm{U}(1)_{e}$ symmetry of electric charge conservation decouples from the dynamics at the critical point, it still has topological contribution. First, odd $N_{mf}$ can be understood as the insertion of a $2\pi$-flux of $\mathrm{U}(1)_{e}$ on top of even $N_{mf}$, i. e., $q$ is increased by $1/2$. Second, the $2\pi$-flux insertion is bound to a $\mathbb{Z}_{2}$ gauge charge. The binding of flux and gauge charge can be seen most evidently from the mutual Chern-Simons $(1/2\pi)\,a\,\mathrm{d}A$ coupling of the the electromagnetic field $A$ and the dynamical gauge field $a$ (cf. Appendix B). Here $a$ is a $\mathrm{U}(1)$ dynamical gauge field which is Higgsed to $\mathbb{Z}_{2}$. [^9] An equivalent way to see this Higgs mechanism is to impose $\mathrm{d}a=0$ and $\oint_{\mathscr{C}}a=0$ or $\pi$, where $\mathscr{C}$ is a closed loop. Its equation of motion asserts
$Q_{a}=\frac{1}{2\pi}\Phi_{A}$
@@
i. e., upon the insertion of $2\pi$-flux of $A$, a $\mathbb{Z}_{2}$ gauge charge is also inserted together with an electric charge. Due to the topological nature, the attached Wilson line preserves the $\mathrm{SO}(3)$ sphere rotation symmetry, which is different from the usual case of a non-topological conformal line. The defect endpoint operators are the $\mathbb{Z}_{2}$-odd part of the Ising CFT, and they can be observed from the eigenstates through the state-operator correspondence on fuzzy sphere, as we will discuss below.
@@
Figure 9: (a) The rescaled bosonic energy gap $\Delta E_{b}N_{mf}^{1/2}$ as a function of $\mu$ in vicinity of the Ising transition for both even and odd $N_{mf}$ calculated at $t=0.1$. The lines with even and odd $N_{mf}$ intersect respectively at $\mu_{c2}\approx 2.0$ except for the smallest system sizes $N_{mf}=6,7$. (b) The rescaled energy spectrum at conformal point $t=0.1,\mu=2.29$ calculated at $N_{mf}=13$ described by the Wilson line endpoint operators in the Ising CFT. We identify some multiplets of the lowest primaries in the spectra and the bars denote their expected value in the CFT. The energy spectrum is rescaled such that $\Delta_{\sigma}=0.518$ and $\Delta_{\partial^{\mu}\sigma}=1.518$.

### v3

@@ 4 Critical Theories
In this section, we discuss the critical theories that describe the transition between the fIQH, MQH, and bPf phases.
The transition from the fIQH to the MQH phase is described by a single free Majorana fermion. One way to see this is that the chiral central charge $c_{-}$ increases by $1/2$ across the transition. This transition is captured by the Lagrangian
@@
where $\chi$ is a (two-component) Majorana fermion, $\bar{\chi}=\chi^{\mathrm{T}}\epsilon$, $\epsilon=i\gamma^{2}$, and $\not{\partial}=\gamma^{\mu}\partial_{\mu}$. The gravitational Chern-Simons term is defined through extending into a 4D manifold $X$ with the original 3D manifold its boundary $M=\partial X$ [80]
@@
where $R$ is the curvature 2-form. A background $\text{CS}_{g}$ level-1 from the background integer quantum Hall state has been included. Adding a positive/negative mass gaps out the Majorana fermion and adds an extra gravitational Chern-Simons response of level $\pm 1/4$.
@@
They recover the chiral central charge of the fIQH and MQH phases. This transition has another conjectural dual description of $\mathrm{SO}(3)_{1}$ coupled to a single free scalar [81, 82, 83]. The condensation of the scalar Higgses the gauge theory to $\mathrm{SO}(2)_{1}$ describing the fIQH phase, and gapping the scalar results in $\mathrm{SO}(3)_{1}$ corresponding to the MQH phase.
The transition from the MQH to the bPf phase falls into a gauged Ising (Ising) universality, i. e. Ising type transition coupled to a $\mathbb{Z}_{2}$ gauge field, or the confinement transition of a $\mathbb{Z}_{2}$ gauge field. One way to understand it is to recognise that bPf can be described by MQH coupled to a $\mathbb{Z}_{2}$ gauge field, so its transition to MQH can be realised by Higgsing the $\mathbb{Z}_{2}$ gauge field. We can also use the language of CS theory to understand the emergence of gauged Ising transition. The MQH and the bPf phase can be described by $\mathrm{SO}(3)_{1}\cong\mathrm{SU}(2)_{2}/\mathbb{Z}_{2}$ and $\mathrm{SU}(2)_{2}$ respectively, and the difference is only a $\mathbb{Z}_{2}$ gauge field. So Higgsing the $\mathbb{Z}_{2}$ centre of the $\mathrm{SU}(2)$ gauge group drives a transition from the bPf phase to the MQH phase. In Appendix B, we discuss the Lagrangian of this confinement transition. This transition can be equivalently considered in terms of anyon condensation. In the bPf phase, although the $\psi$ anyon is fermionic, combined with the local fermion $f$, $f^{\dagger}\psi$ is bosonic with statistical angle $\theta_{f^{\dagger}\psi}=0$. Across the transition, this composite particle $\langle f^{\dagger}\psi\rangle\neq 0$ is condensed. Due to the braiding of the anyons $\psi$ and $\sigma$, $\sigma$ will also be confined, resulting in the MQH with trivial topological order. Since this transition only changes the neutral sector, the chiral central charge $c_{-}=3/2$ remains the same across the transition. This MQH-bPf transition can also be described as the adjoint QCD of $\mathrm{SU}(2)_{1}$ Chern-Simons theory coupled to a fermion in the adjoint representation of the gauge group [84]. When the fermion is gapped, it contributes a Chern-Simons level $\pm 1$ depending on the sign of the mass. This drives a transition between $\mathrm{SU}(2)_{2}$ topological order and a trivial topological order $\mathrm{SU}(2)_{0}$, which only differs from the MQH phase by an invertible TQFT. This transition is conjectured to be dual to $\mathrm{O}(1)_{-3}$ coupled to a real scalar through the level-rank duality, which is equivalent to the Ising universality.
@@
Figure 4: The rescaled energy spectrum at the conformal points (a) $t=0.3,\mu=0.0$ described by the free-Majorana-fermion CFT and (b) $t=0.1,\mu=2.25$ described by the Ising CFT. We identify some multiplets of the lowest primaries in the spectra and the bars denote their expected value in the CFT. The spectra are calibrated by $\Delta_{T}=3$. The brown circles denote states not identified. In the calculation, we fix $N_{mf}=12$.
At the fIQH-MQH transition, we observe that most of the operators are close to integer levels and they can be classified into multiplets of the lowest primary operators in the free-Majorana-fermion CFT (Figure 4a), viz. the fermion $\chi$ at $\Delta=1,l=1/2$, the mass field $\bar{\chi}\chi=\epsilon_{\alpha\beta}\chi^{\alpha}\chi^{\beta}$ at $\Delta=2,l=0$, the conserved stress tensor $T^{\mu\nu}=\tfrac{i}{2}\bar{\chi}_{a}\alpha(\gamma^{(\mu})^{\alpha}{}_{\beta}\partial^{\nu)}\chi^{\beta}$ at $\Delta=3,l=2$, $\bar{\chi}\chi\partial\chi$ at $\Delta=4,l=3/2$, etc. We highlight that spectrum contains $l\in\mathbb{Z}+\tfrac{1}{2}$ states corresponding to fermionic operators. The free-Majorana-fermion CFT will be analysed in more detail in Section 5.
At the MQH-bPf transition, although there are many non-conformal gapped excitations at relatively low energy at small system size, the low lying states for each integer $l\leq 2$ can be identified as operators in the Ising CFT (Figure 4b). The Ising CFT consists all the $\mathbb{Z}_{2}$-even operators in the Ising CFT. We can identify $\epsilon$ at $\Delta=1.412$ and its several descendants, the stress tensor $T^{\mu\nu}$ and $\epsilon^{\prime}$ at $\Delta=3.830$ in the spectrum. The Ising CFT will be discussed in more detail in Section 6.
5 Free-Majorana-Fermion CFT
In this section, we focus on the free-Majorana-fermion CFT on the fuzzy sphere. In this CFT, the Majorana fermion $\chi$ has scaling dimension $\Delta=1$ and $\mathrm{SU}(2)$ spin $l=1/2$. The spectrum of scaling operator, including the primaries and descendants, can be determined with the help of harmonic oscillators [35]. Here we list the scaling dimension $\Delta$ and spin $l$ of the low lying operators:
@@ 5 Free-Majorana-Fermion CFT
Figure 5: The spectrum with $\Delta\lesssim 4$ for various systems sizes in the free-Majorana-fermion CFT. The scaling dimensions are calibrated by the stress tensor with $\Delta_{T}=3$. The grey bars denote the theoretical value for comparison. In the calculation, we fix $t=0.3$ and $\mu=0$. The verticle gridlines with interval $0.5$ separates differnt spins, which are either integers or half-integers.
We find that the low lying spectrum agrees well with the theory (Figure 5). All the states up to $\Delta=4$ are CFT states, and all the theoretically expected operators appear in the spectrum close to integer scaling dimensions. The discrepancy is within $2.5\%$ for all the $\Delta\leq 3$ operators and most $\Delta=4$ operators for our maximal accessible size $N_{mf}=13$. This precision is remarkable with almost no fine-tuning at such small size even compared with similar fuzzy-sphere model (e. g. the free real scalar [35]). With increasing system size, the scaling dimension gradually approaches to expected integer value (Figure 6a–d). We quantitatively evaluate this through the cost function consisting of all the operators with $\Delta\leq\Delta_{c}=2,3,4$. We find these cost functions all scales to zero as $Q\sim N_{mf}^{-\alpha/2}$ where $\alpha=3.4,3.6,3.2$ respectively for $\Delta_{c}=2,3,4$ (Figure 6e).
Figure 6: (a–d) The finite-size scaling of scaling dimensions of various operators in the free-Majorana-fermion CFT. The grey lines denote the theoretical value for comparison. (e) The finite-size scaling of cost function consisting of operators with $\Delta\leq\Delta_{c}=2,3,4$. In the calculation, we fix $t=0.3$ and $\mu=0$.
@@
$\eta(\mathbf{r})$ $=f^{\dagger}(\mathbf{r})b(\mathbf{r})=\alpha\chi^{(1/2)}(\mathbf{r})+\dots$
$\eta^{\dagger}(\mathbf{r})$ $=b^{\dagger}(\mathbf{r})f(\mathbf{r})=\bar{\alpha}\chi^{(-1/2)}(\mathbf{r})+\dots,$ (8)
@@
$\langle\chi^{\alpha}(x_{1})\bar{\chi}_{\beta}(x_{2})\rangle=\frac{(\gamma_{\mu})^{\alpha}{}_{\beta}x_{12}^{\mu}}{(x_{12}^{2})^{\Delta_{\chi}+1/2}},$ (9)
where $\Delta_{\chi}=1$, $x_{1,2}\in\mathbb{R}^{3}$, $x_{12}=x_{1}-x_{2}$, $\alpha,\beta=\pm$ are spinor indices, and $\gamma$ are the Pauli matrices. Taking the polarisations $\chi^{(s)}(\mathbf{r})$ on the sphere, their two-point function
$\langle 0|\chi^{\dagger\,(s_{1})}(\mathbf{r}_{1})\chi^{(s_{2})}(\mathbf{r}_{2})|0\rangle=\frac{2(-1)^{s_{1}}\delta_{s_{1}s_{2}}}{R^{2\Delta_{\chi}}(2\sin\frac{1}{2}\theta_{12})^{2\Delta_{\chi}}}$ (10)
has the same form as scalars on the sphere, where $\theta_{12}$ is the angular distance between the two points.
Figure 7: (a) The dimensionless correlation function of the Majorana fermion $\chi$ Eq. (11) at diffrent system sizes. The grey dashed line denotes $(2\sin\theta_{12}/2)^{-2}$ required by conformal symmetry. (b) The scaling dimension $\Delta_{\chi}$ extractd from the corrlation function through different ways in Eqs. (12) as a function of system size. In the calculation, we fix $t=0.3$ and $\mu=0$.
@@
$C_{\chi}^{\ast}(\theta_{12})=\tfrac{1}{2i}R^{2\Delta_{\chi}}\langle 0|\chi^{\dagger\,(1/2)}(\mathbf{r}_{1})\chi^{(1/2)}(\mathbf{r}_{2})|0\rangle.$ (11)
whose exact expression is given in Eq. (53), and plot the results in Figure 7a. As the system size increases, the numerical result gradually approaches its theoretical value $(2\sin\theta_{12}/2)^{-2\Delta_{\chi}}$. From this correlation function, the scaling dimension $\Delta_{\chi}$ can be extracted in several ways
$\left.C_{\chi}^{\ast}(\theta_{12})\right|_{\theta_{12}=\pi}$ $=2^{-2\Delta_{\chi}}$ (12a)
$\left.\frac{(C^{\ast}_{\chi})^{\prime\prime}(\theta_{12})}{C^{\ast}_{\chi}(\theta_{12})}\right|_{\theta_{12}=\pi}$ $=\frac{\Delta_{\chi}}{2}$ (12b)
$\int_{-1}^{1}\mathrm{d}\cos\theta_{12}\,\sin\frac{\theta_{12}}{2}C^{\ast}_{\chi}(\theta_{12})$ $=\frac{2^{2(1-\Delta_{\chi})}}{3-2\Delta_{\chi}}.$ (12c)
@@ 6 Gauged Ising CFT
Numerically, we can indeed identify $\epsilon,\epsilon^{\prime},T^{\mu\nu}$ and their descendants up to $\Delta<5$ and $l\leq 2$ in the spectrum (Figure 8a). In the calculation, we fix $t=0.1$ with a relatively optimal conformal symmetry, and choose $\mu$ for each system size by optimising a cost function consisting of $\epsilon,\partial^{\mu}\epsilon,\partial^{\mu}\partial^{\nu}\epsilon,\Box\epsilon$ and $T^{\mu\nu}$. The optimal $\mu$ for each size is listed in Table 1 (the meaning of the right panel will be explained shortly). We note that although there are non-conformal states due to the lowest gapped excitations in the proximate bosonic Pfaffian phase, the low energy spectrum with $l\leq 2$ show relatively clear conformal signature. At large enough system size, the low energy specctrum should be dominated by CFT states and these gapped excitations will move higher, but at the sizes available, these excitations appear at relatively low energy at large $l$. As the bosonic Pfaffian state is only compatible with even $N_{mf}$, and the optimal conformal point drifts with system size, we leave the finite-size scaling analysis to future work.
@@
Figure 8: The spectrum of the identified CFT states with $\Delta<5$ for various systems sizes. The grey bars denote the theoretical value for comparison. (a) The bulk spectrum at even $N_{mf}$ corresponding to the $\mathbb{Z}_{2}$-even part of the Ising CFT spectrum. The scaling dimensions are calibrated by the stress tensor with $\Delta_{T}=3$. (b) The Wilson-line end-point spectrum at odd $N_{mf}$ corresponding to the $\mathbb{Z}_{2}$-odd part of the Ising CFT spectrum. The scaling dimensions are calibrated by setting $\Delta_{\sigma}=0.518$ and $\Delta_{\partial^{\mu}\sigma}=1.518$. In the calculation, we fix $t=0.1$ and choose $\mu$ as in Table 1.
@@
On the other hand, the $\mathbb{Z}_{2}$-odd operators can be made gauge invariant by attaching a Wilson line (of $\mathbb{Z}_{2}$ gauge field) to it. In other words, the $\mathbb{Z}_{2}$-odd operators become end-point operators of a line defect in the gauged Ising CFT. Although the defect end-point operators arise generically in conformal line defects and have been discussed in various contexts [85, 86, 49, 87], in the setting of $\mathbb{Z}_{2}$ (or any discrete) gauge theory, the Wilson line defect is not only conformal but also topological. In the bPf-MQH transition, we can straightforwardly realise the topological Wilson line on the fuzzy sphere by leveraging the even-odd effect of $N_{mf}$.
@@
We first note that the bosonic Pfaffian state is sensitive to the parity of particle number, and is only compatible with even $N_{mf}$. The choice of the odd $N_{mf}$, on the other hand, is equivalent to the insertion of a gauge charge at the centre of the sphere as the end-point of a Wilson line defect. The reason is that although the $\mathrm{U}(1)_{e}$ symmetry of electric charge conservation decouples from the dynamics at the critical point, it still has topological contribution. First, odd $N_{mf}$ can be understood as the insertion of a $2\pi$-flux of $\mathrm{U}(1)_{e}$ on top of even $N_{mf}$, i. e., $q$ is increased by $1/2$. Second, the $2\pi$-flux insertion is bound to a $\mathbb{Z}_{2}$ gauge charge. The binding of flux and gauge charge can be seen most evidently from the mutual Chern-Simons $(1/2\pi)\,a\,\mathrm{d}A$ coupling of the the electromagnetic field $A$ and the dynamical gauge field $a$ (cf. Appendix B). Here $a$ is a $\mathrm{U}(1)$ dynamical gauge field which is Higgsed to $\mathbb{Z}_{2}$. [^11] An equivalent way to see this Higgs mechanism is to impose $\mathrm{d}a=0$ and $\oint_{\mathscr{C}}a=0$ or $\pi$, where $\mathscr{C}$ is a closed loop. Its equation of motion asserts
$Q_{a}=\frac{1}{2\pi}\Phi_{A},$
@@
i. e., upon the insertion of $2\pi$-flux of $A$, a $\mathbb{Z}_{2}$ gauge charge is also inserted together with an electric charge. Due to the topological nature, the attached Wilson line preserves the $\mathrm{SO}(3)$ sphere rotation symmetry, which is different from the usual case of a non-topological conformal line. The defect end-point operators are the $\mathbb{Z}_{2}$-odd part of the Ising CFT, and they can be observed from the eigenstates through the state-operator correspondence on the fuzzy sphere, as we will discuss below.
@@
Figure 9: (a) The rescaled bosonic energy gap $\Delta E_{b}N_{mf}^{1/2}$ as a function of $\mu$ in vicinity of the Ising transition for both even and odd $N_{mf}$ calculated at $t=0.1$. The lines with even and odd $N_{mf}$ intersect respectively at $\mu_{c2}\approx 2.0$ except for the smallest system sizes $N_{mf}=6,7$. (b) The rescaled energy spectrum at conformal point $t=0.1,\mu=2.29$ calculated at $N_{mf}=13$ described by the Wilson-line end-point operators in the Ising CFT. We identify some multiplets of the lowest primaries in the spectra and the bars denote their expected value in the CFT. The energy spectrum is rescaled such that $\Delta_{\sigma}=0.518$ and $\Delta_{\partial^{\mu}\sigma}=1.518$.